# Explicit surjectivity of Galois representations of products of elliptic curves over function fields

### v2

@@ 1. Introduction
Let $K$ be the function field of a smooth, projective, and geometrically integral curve $C$ over a perfect field $k$. Let $E_{1},\ldots,E_{n}$ be pairwise non-isogenous elliptic curves over $K$ with $j(E_{i})\not\in\overline{k}$ for each $1\leq i\leq n$, and let
@@
$c(g)$ $:=2+\max\left\{\ell\mid\dfrac{q-(6+3e_{2}(q)+4e_{3}(q))}{12}\leq g\right\},$
$C(g)$ $=3176523\cdot\max\{1,g(K)^{3/2}\},$
@@ 2.2.2. Separability and biseparability
is also an isogeny of degree $p^{\textnormal{dim}(X)}$. We have that $V_{X}\circ F_{X}=[p]_{X}$ and $F_{X}\circ V_{X}=[p]_{X^{(p)}}$, and these homomorphisms are Cartier dual to one-another [EvdGM, Prop 5.19, Prop 5.20]. The separability degree of $V_{X}$ is intimately related to the structure of the $p$-power torsion of $X$.
@@ 3. An isogeny degree bound
This follows from the same argument given for [GP22, Proposition 6.7], but using the cruder degree bound of [GP22, Proposition 6.6] in place of [GP22, Prop 6.5]. ∎
@@
The minimality of the degree of $\varphi_{0}$ forces $\varphi_{0}$ to be cyclic. Supposing otherwise, there exist integers $m_{1},m_{2},m_{3},m_{4}\in\mathbb{Z}^{+}$ with $m_{1}\mid m_{i}$ for all $1\leq i\leq 4$ such that
$\textnormal{ker}(\varphi_{0})(\overline{K})\simeq\mathbb{Z}/m_{1}\mathbb{Z}\times\mathbb{Z}/m_{2}\mathbb{Z}\times\mathbb{Z}/m_{3}\mathbb{Z}\times\mathbb{Z}/m_{4}\mathbb{Z},$
and it follows that $\varphi_{0}$ must factor through $m_{1}]:E_{1}\times E_{2}\rightarrow E_{1}\times E_{2}$. If $m_{1}>1$, then by Proposition 2.4 there exists an isogeny $\psi:E_{1}\times E_{2}\rightarrow X$ with $\varphi_{0}=\psi\circ[m_{1}]$. By comparing degrees, we find that $\psi$ is both biseparable and of degree strictly less than $d_{0}$, contradicting the minimality of the degree of $\varphi_{0}$. Therefore, $m_{1}=1$, and repeated application of this argument gives $m_{2}=m_{3}=1$ as well, such that $\varphi_{0}$ is cyclic.
@@
Note that $\textnormal{ker}(\varphi_{1})\times\{0_{2}\}\leq\textnormal{ker}(\varphi_{0})(\overline{K})$ and $\{0_{1}\}\times\textnormal{ker}(\varphi_{2})\leq\textnormal{ker}(\varphi_{0})(\overline{K})$. Moreover, if $(P,Q)\in\textnormal{ker}(\varphi_{0})(\overline{K})$, then with $0_{X}$ denoting the identity of $X$ we find
@@
Now, by the same arguments as in the proof of [GP22, Prop 6.7] and the fact that $d_{1},d_{2}\mid d_{0}$, both $\varphi_{1}$ and $\varphi_{2}$ are cyclic and biseparable with kernel stable under the action of $G_{K}$. Thus $\varphi_{i}$ induces a $K$-rational point on $X_{0}(d_{i})$ for $i=1,2$. Lemma 3.1 then provides
@@ 4. Effective surjectivity for products
An effective result in the case of a single elliptic curve was obtained by Cojocaru–Hall [CH05, Thmeorem 1.1], and we will make use of their theorem here. To state it, we recall the following setup. For $E$ an elliptic curve over $K$ and $\ell$ a rational prime not equal to $\textnormal{char}(K)$, let
@@
$c(g):=2+\max\left\{q\mid\dfrac{q-(6+3e_{2}(q)+4e_{3}(q))}{12}\leq g\right\}.$
Then for all $\ell\geq c(g)$, we have $H_{\ell}=\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$.
@@
To state our theorem, we recall the following setup from Section 1. Let $E_{1},\ldots,E_{n}$ with $n\geq 2$ be non-isotrivial, pairwise non-isogenous elliptic curves over a function field $K$, and let
@@ 5. An average result over families of products
From the Siegel–Walfisz theorem (see [IK04, Cor. 5.29]), we have that if $\ell\ll\left(\log{x}\right)^{B}$ for some real number $B>0$ then $\pi(x;\ell,\delta)\gg\frac{x}{\ell\log{x}}$. Thus, for $\ell$ and $t_{0}$ as above we have

### v3

@@ 1. Introduction
Let $K$ be the function field of a smooth, projective, and geometrically integral curve $C$ of genus $g$ over a perfect field $k$. Let $E_{1},\ldots,E_{n}$ be pairwise non-isogenous elliptic curves over $K$ with $j(E_{i})\not\in\overline{k}$ for each $1\leq i\leq n$, and let
@@
$c(g)$ $:=2+\max\left\{q\textnormal{ prime }\bigm|\dfrac{q-(6+3e_{2}(q)+4e_{3}(q))}{12}\leq g\right\},$
$C(g)$ $=3176523\cdot\max\{1,g^{3/2}\},$
@@ 2.2.2. Separability and biseparability
is also an isogeny of degree $p^{\textnormal{dim}(X)}$. We have that $V_{X}\circ F_{X}=[p]_{X}$ and $F_{X}\circ V_{X}=[p]_{X^{(p)}}$, and these homomorphisms are Cartier dual to one-another [EvdGM, Proposition 5.19, Proposition 5.20]. The separability degree of $V_{X}$ is intimately related to the structure of the $p$-power torsion of $X$.
@@ 3. An isogeny degree bound
This follows from the same argument given for [GP22, Proposition 6.7], but using the cruder degree bound of [GP22, Proposition 6.6] in place of [GP22, Proposition 6.5]. ∎
@@
Note that $\textnormal{ker}(\varphi_{1})(\overline{K})\times\{0_{2}\}\leq\textnormal{ker}(\varphi_{0})(\overline{K})$ and $\{0_{1}\}\times\textnormal{ker}(\varphi_{2})(\overline{K})\leq\textnormal{ker}(\varphi_{0})(\overline{K})$. Moreover, if $(P,Q)\in\textnormal{ker}(\varphi_{0})(\overline{K})$, then with $0_{X}$ denoting the identity of $X$ we find
@@
We claim that both $\varphi_{1}$ and $\varphi_{2}$ are cyclic isogenies of elliptic curves. Suppose, for example, that $\varphi_{1}$ is not cyclic, so there exists some $m\mid d_{1}$ with $m>1$ such that $\varphi_{1}$ factors through the endomorphism $m]_{E_{1}}:E_{1}\to E_{1}$. It follows that $\varphi_{0}$ factors through the endomorphism
$\eta:=[m]_{E_{1}}\times\textnormal{Id}_{E_{2}}\in\operatorname{End}(E_{1}\times E_{2})\cong\mathbb{Z}\times\mathbb{Z},$
and hence by Proposition 2.4 there exists an isogeny $\psi:E_{1}\times E_{2}\rightarrow X$ with $\varphi_{0}=\psi\circ\eta$. By comparing degrees, we find that $\psi$ is both biseparable and of degree strictly less than $d_{0}$, contradicting the minimality of the degree of $\varphi_{0}$.
Now, by the same arguments as in the proof of [GP22, Proposition 6.7] and the fact that $d_{1},d_{2}\mid d_{0}$, both $\varphi_{1}$ and $\varphi_{2}$ are cyclic and biseparable with kernel stable under the action of $G_{K}$. Thus $\varphi_{i}$ induces a $K$-rational point on $X_{0}(d_{i})$ for $i=1,2$. Lemma 3.1 then provides
@@ 4. Effective surjectivity for products
An effective result in the case of a single elliptic curve was obtained by Cojocaru–Hall [CH05, Theorem 1.1], and we will make use of their theorem here. To state it, we recall the following setup. For $E$ an elliptic curve over $K$ and $\ell$ a rational prime not equal to $\textnormal{char}(K)$, let
@@
$c(g):=2+\max\left\{q\textnormal{ prime }\bigm|\dfrac{q-(6+3e_{2}(q)+4e_{3}(q))}{12}\leq g\right\}.$
We then have for all primes $\ell\geq c(g)$ with $\ell\neq\textnormal{char}(K)$ that $H_{\ell}=\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$.
@@
To state our theorem, we recall the following setup from Section 1. Let $E_{1},\ldots,E_{n}$ with $n\geq 2$ be non-isotrivial, pairwise non-isogenous elliptic curves over a function field $K$ of genus $g$, and let
@@ 5. An average result over families of products
From the Siegel–Walfisz theorem (see [IK04, Corollary 5.29]), we have that if $\ell\ll\left(\log{x}\right)^{B}$ for some real number $B>0$ then $\pi(x;\ell,\delta)\gg\frac{x}{\ell\log{x}}$. Thus, for $\ell$ and $t_{0}$ as above we have